OpenStudy (anonymous):

HELP PLEASEEE !!! DOES ANYBODY KNOW HOW TO SOLVE THE SYSTEM OF EQUATIONS THROUGH elimination x+4y=10 -x+y=4

OpenStudy (shamil98):

Elimination. Just add the two equations together like so. x+4y = 10 +(-x) + y = 4 ------------- Can you do this first step and tell me what your resulting equation is?

OpenStudy (anonymous):

isnt it 5y =14 ?

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

Yes. 5y = 14 Now, just solve for y.

OpenStudy (anonymous):

y=14/5